OpenStudy (anonymous):

There are 3 black and 7 white balls in the box, identical in touch. You draw 3 balls randomly. If you find 1 black ball among the drawn ones, you will win $1, if you find 2 black balls, you win $2 and you will win $10, if all 3 are black. What are the chances to win at least 1 buck?

OpenStudy (anonymous):

Pr(win at least $1) = Pr(draw at least 1 black ball) = 1 - Pr(draw no black balls) = 1 - 7/10 x 6/9 x 5/8